A test procedure requires 500mL of a 4N phosphoric acid solution. The concentrated phosphoric acid available has a specific gravity of 1.69 and an assay of 85%. How would you prepare this solution? (MW H3PO4 = 98) (Valence is 3)
Answer #1

We know that

normality = Molarity * valence

Given that

normality of the solution = 4 N

So, molarity of the solution = normality /valence = 4/3 = 1.333333 M

Volume of 4 N solution required = 500 mL = 0.5 L

Number of moles of H3PO4 = M*V = 1.333333 M * 0.5 L = 0.666667 mol

Mass of H3PO4 = moles * mol.wt. = 0.666667 mol * 98 g/mol = 65.33333 g

Given that

Mass percent = 85 %

Mass percent = mass of solute/mass of the solution *100

85 = 65.33333/mass of the solution * 100

Mass of the solution = 65.33333/85*100 = 76.86 g

Given that

specific gravity = 1.69

We know that

specific gravity = density of the solution/density of water

density of the solution = specific gravity * density of water

density of the solution = 1.69 * 1.0 g/mL = 1.69 g/mL

Volume of the solution = mass/density = 76.86 g/(1.69 g/mL) = 45.48 mL

Volume of water to be added = 500 – 45.48 = 454.52 mL

Procedure :

Take 454.52 mL of water into a 500 mL volumetric flask and then add slowly 45.48 mL of concentrated Phosphoric acid